OpenStudy (anonymous):

A local high school offers a total of 8 language classes, some Spanish and some French classes. If the number of Spanish classes offered is 2 more than twice the number of French classes, how many Spanish classes are actually offered

OpenStudy (anonymous):

call x, the number of french classes. 2x+2+x=8 3x=6 x=2 so the number of spanish classes offered must be 8-2=6

OpenStudy (anonymous):

s=2+2f s+f=8 2+2f+f=8 3f=6 f=2 s=6

OpenStudy (anonymous):

s=Spanish classes f=French classes

OpenStudy (anonymous):

Your Both Correct nice Work Good Explaining ;)